OpenStudy (anonymous):

Name the property the equation illustrates (ab)3 = a(b3) A. Commutative property of addition B. Associative property of Multiplication C. Associative Property of Addition D. Inverse Property of Multiplication

OpenStudy (notamathgenius):

It's B or D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Right?

OpenStudy (notamathgenius):

You're right, inverse is just taking the formula and flipping it around.

OpenStudy (anonymous):

Was wrong.

OpenStudy (notamathgenius):

WHAT O_O It's right I just did this

OpenStudy (anonymous):

Yeah I coulda swore it was right to.

OpenStudy (notamathgenius):

They took 2 "numbers" and flipped them around the formula to make INVERSE that's not good D: I'm sorry

OpenStudy (anonymous):

What about this one? 9 + (-9) = 0

OpenStudy (anonymous):

o_O?

OpenStudy (notamathgenius):

A negative times a positive is a negative xD sorry

OpenStudy (anonymous):

^Additive Inverse

OpenStudy (anonymous):

Right. Addition property of 0?

OpenStudy (anonymous):

Oh..

OpenStudy (anonymous):

Yeah that was right. :p

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Additive_inverse

OpenStudy (anonymous):

This one has to be inverse

OpenStudy (anonymous):

8 + 3.4 = 3.4 + 8

OpenStudy (anonymous):

Commutative Property of Addition

OpenStudy (anonymous):

Whaat.-_- I suck at this, lmao.

OpenStudy (anonymous):

This ones confusing.

OpenStudy (anonymous):